OpenStudy (anonymous):

please help let the function f be defined by f(x)=x^2-2x+1 for all numbers x. if f(a)=4, which of the following could be the value of a? a) -1 b) 0 c) 1 d)2 e)4

OpenStudy (aum):

f(x) = x^2 - 2x + 1 f(a) = a^2 - 2a + 1 = 4 a^2 - 2a + 1 - 4 = 0 a^2 - 2a - 3 = 0 a^2 - 3a + a - 3 = 0 a(a - 3) + 1(a - 3) = 0 (a-3)(a+1) = 0 a = 3, -1

OpenStudy (anonymous):

thank you so much